### v1

# Modelling the 5G Energy Consumption using Real-world Data: Energy Fingerprint is All You Need

Abstract
The introduction of fifth-generation (5G) radio technology has revolutionized communications, bringing unprecedented automation, capacity, connectivity, and ultra-fast, reliable communications. However, this technological leap comes with a substantial increase in energy consumption, presenting a significant challenge. To improve the energy efficiency of 5G networks, it is imperative to develop sophisticated models that accurately reflect the influence of base station (BS) attributes and operational conditions on energy usage. Importantly, addressing the complexity and interdependencies of these diverse features is particularly challenging, both in terms of data processing and model architecture design.
This paper proposes a novel 5G base stations energy consumption modelling method by learning from a real-world dataset used in the ITU 5G Base Station Energy Consumption Modelling Challenge in which our model ranked second. Unlike existing methods that omit the Base Station Identifier (BSID) information and thus fail to capture the unique energy fingerprint in different base stations, we incorporate the BSID into the input features and encoding it with an embedding layer for precise representation. Additionally, we introduce a novel masked training method alongside an attention mechanism to further boost the model’s generalization capabilities and accuracy. After evaluation, our method demonstrates significant improvements over existing models, reducing Mean Absolute Percentage Error (MAPE) from 12.75% to 4.98%, leading to a performance gain of more than 60
Index Terms:
5G, Base Station, Energy Consumption, Deep Learning
I Introduction
The advent of the fifth-generation (5G) network introduces advanced capabilities that significantly benefit various applications, including virtual reality, vehicular networks and the Internet of Things (IoT). However, this technological leap comes with a substantial increase in energy consumption. Compared to its predecessor, the fourth-generation (4G) network, the energy consumption of the 5G network is approximately three times higher [1]. Notably, energy costs constitute about 22.5% of the total expenditure for mobile network operators [2], a figure that is on the rise. A critical factor in this context is the radio access network (RAN), particularly the base stations (BSs), which account for over 70% of the network’s energy usage [3].
Commercial networks feature BSs with various architectures, such as remote radio units or antenna processing units. The power consumption profile of these BSs, which are equipped with diverse capabilities such as varying numbers of carriers and antennas, are further influenced by their configuration settings, impacting their energy consumption. Despite their substantial energy requirements, there exist methods to mitigate RAN consumption, primarily by disabling specific hardware components within the BSs. These energy-saving methods can be activated manually or automatically under certain traffic conditions. Understanding the interplay of these factors is essential for optimizing future RAN energy efficiency.
The precise assessment of BS energy consumption is pivotal in the modelling of energy consumption in 5G network, a topic that has been a longstanding challenge and has consistently garnered attention, with numerous studies conducted to investigate this issue. The work presented in [4] introduced a popular analytical power consumption model highlighting the linear correlation between the power output of the BS and its power usage. This model is further extended in [5], incorporating considerations for the massive multiple-input multiple-output (mMIMO) architecture and multiple energy-efficient techniques. Subsequent work in [6] expanded the model in [4] considering a linear increase of the power consumption with the number of mMIMO transceivers. A comprehensive and detailed framework for understanding power consumption was proposed in [7], taking into account factors like the mMIMO architecture, downlink and uplink phases, and the number of users multiplexed.
So far, several emerging tasks such as sensing [8], mobile edge computing [9], edge learning and inference [10, 11] have been applied to extend BS applications for next-generation networks. These additional tasks not only increase the BSs’ energy consumption but also introduce new characteristics relevant to energy consumption modelling, directly impacting how energy efficiency is understood and optimized. The increasing complexity of mobile networks, coupled with the diverse capabilities of BSs, necessitates the use of machine learning models that can effectively account for the impact of various parameters on energy consumption. Most recently, SRCON (Simulated Reality of Communication Networks) has emerged as a revolutionary paradigm in network simulation. Detailed in  [12, 13], SRCON leverages an innovative mix of white-box and black-box models to mimic the stochastic behaviors observed in real-world 4G/5G mobile networks, offering a robust framework for accurate simulation. Within this framework, [14, 15] develops a machine learning model, trained on extensive network measurement data, which significantly improves the accuracy of energy consumption estimations compared to traditional mathematical modelling.
However, existing modelling methods do not take the complexity and heterogeneity of real commercial products into consideration. Such complexity is reflected in various previously unconsidered features, such as hardware characteristics and physical environments, which we refer to as the energy fingerprint. Omitting the said energy fingerprint may lead to the incomplete input feature space and, what’s worse, the learnt model may suffer from severe one-to-many issue, i.e., the same input features may be mapped to different prediction outcomes, posing significant challenges to achieving accurate prediction outcomes.
To alleviate the one-to-many issue, we propose a novel modelling method based on the real-world dataset from the ITU 5G Base Station Energy Consumption Modelling Challenge [16] where our solution ranked second among more than two hundred teams throughout the world. Specifically, our model incorporates the Base Station Identifier (BSID) feature as input, which was ignored by previous methods, to capture the inherent energy fingerprint of different BSs. Furthermore, we utilize an embedding layer for BSID feature. This method allows a more precise representation of the BSID feature and addresses the shortcomings of the classic One-Hot encoding, which leads to a prohibitively high-dimension of the input feature space and thus an unnecessarily large model size of the neural network to accommodate the high-dimension input, hindering the model’s ability to generalize accurately. More importantly, it is found that the embedding layer can project the BSID into a latent space in which BSs are clustered according to their energy fingerprint, justifying the capability of the proposed model in energy fingerprint extraction. Moreover, to endow the embedding layer with generalization capability to handle the unseen BSID in the test data, we further introduce a new masked training technique to boost the model’s generalization ability, making it more applicable in real-world settings. Last but not least, we design the attention mechanism for further automatic feature selection besides the manual selection before training. Extensive experiments on real-world 5G network measurement data are conducted to examine the superiority of the proposed design over the existing baselines in diverse test scenarios. The Mean Absolute Percentage Error (MAPE) improved from 12.75% in the existing method to 4.98%, resulting in a performance enhancement of over 60% for the algorithm.
II 5G Radio Unit Dataset
In this section, we introduce the 5G Radio Unit Dataset. The dataset used in our study comprised 102,705 hourly measurements collected over a period of 8 days. Each sample encapsulates a variety of features, classified into four main categories:
  • Base Station features: These include hardware attributes and configuration details of the BS, which are key in predicting energy consumption. Features in this category are the number of antennas (Antennas), the transmission mode (Mode), the BSID, and the type of radio unit (RUType). BSs with similar hardware attributes and configurations often show similar energy consumption patterns.
  • Cell-Level features: These include hardware attributes and configuration details of each of the cells operated by the BS. Features in this category are Load, activation levels of six different Energy Saving Modes (ESMode1-6), Frequency, Bandwidth, and maximum transmit power (TXpower). These dynamic features are crucial for understanding the hourly energy consumption of the BSs. Additionally, a BS can support up to four cells, with one serving as the primary and the others almost always remaining inactive.
  • Time features: The dataset includes day and hour fields, indicating when the measurement samples were collected.
  • Energy Consumption: The dataset includes hourly measurements of energy consumption.
The dataset is split into training and testing sets to test our model’s generalization capabilities, with certain BSs only present in the testing set. The training set consists of 92,629 samples. The testing set includes 10,076 samples, of which 3,067 are from BSs not encountered during training (unknown BS), and 7,009 are from BSs that appear in the training set (known BS). All BS refers to the entirety of both known and unknown BSs combined. The training data includes 10 distinct RUType categories, while the testing set introduces two additional RUTypes not seen in the training, challenging the model’s ability to adapt to new RUTypes.
III Problem Statement
By analysis of the collected data, there exists a “one-to-many” issue encountered in the datasets, meaning that different energy consumption values are reported for identical input feature values. To better illustrate the said issue, we depict Fig. 1, a scatter plot, in which the labels (i.e., the energy consumption values) of selected data samples are plotted against one of the most important features (i.e., the load of the primary cell as identified in [15]) with other features kept identical.
Three key observations can be made from Fig. 1 as listed below:
  • Energy consumption exhibits a roughly linear relationship with the load of the primary cell.
  • Without distinguishing BSID, as depicted by the gray area in the figure, a common load value corresponds to high variance in energy consumption.
  • When distinguishing by BSID, as depicted by the red, green, and blue areas, a common load value corresponds to much lower variance in energy consumption.
Combining these observations, we can conclude that distinguishing BSID helps a lot in reducing the variance of energy consumption predictions. This gain comes from the fact that incorporating BSID as model input feature enables us to unveil the subtle hardware differences across various BSs, effectively capturing the unique BS energy fingerprint for more accurate predictions.
Figure 1: Illustration of “one-to-many” issue.
IV Deep Learning Model Architecture and Training
In this section, we present our methodology, encompassing the input processing technique, the architecture of our deep learning model, and our innovative masked training approach, aimed at enhancing generalization across different BSs to address the one-to-many issues presented in Section III.
IV-A Input Feature Processing
Our input feature processing is designed to address the diversity of input features through distinct feature encoding methods. Categorical inputs, except for BSID, are processed using One-Hot encoding, which converts them into a binary format, maintaining categorical distinctions. Numerical inputs, on the other hand, are inputted directly into the model in their original form to utilize their inherent magnitude variations. However, for numerical parameters like Antennas, Bandwidth, and Frequency with limited value ranges, we also apply One-Hot encoding to capture subtleties.
For the BSID feature, characterized by its extensive range of classes, an embedding technique is used, transforming the information into a compact numerical vector, thereby balancing dimensionality with information retention.
This tailored approach to input processing has led to improved performance, as detailed in Section IV. The encoding process adopted for all input parameters is outlined in Table I.
TABLE I: Input parameters.
Feature Class Feature Type Feature Encoding
BS-level RUType Categorical One-Hot
Mode Categorical One-Hot
Antennas Numerical One-Hot
BSID Categorical Embedding
Cell-level Load Numerical Direct Input
ESMode1-6 Numerical Direct Input
TXpower Numerical Direct Input
Frequency Numerical One-Hot
Bandwidth Numerical One-Hot
Time Day Categorical One-Hot
Hour Categorical One-Hot
IV-B Deep Learning Model Design
Fig. 2(a) illustrates the architecture of our proposed framework. At its core, the model employs a multilayer perceptron (MLP) structure, and includes specialized modules for embedding and adaptive weight adjustment. The model consists of two hidden layers: the first layer contains 128 neurons, while the second layer comprises 64 neurons. The final layer directly outputs the predictions.
BSID Embedding
The efficient processing of BSID is a critical aspect, especially considering the large number of BSs included in our dataset. The use of traditional One-Hot encoding for BSIDs produces sparse vectors, which can become inefficient.
To tackle this challenge, we have integrated an embedding technique into our model. This approach transforms each categorical BSID into a dense, continuous vector. Specifically, each BSID is represented as an index BSID in a lookup table $\boldsymbol{W}_{\textrm{emb}}$. Initially, each embedding is randomly assigned a compact 64-dimensional numerical vector, $\boldsymbol{x}_{\textrm{BSID}}$. During training, automatic differentiation updates embedding values directly, refining them implicitly without converting indices to One-Hot vectors, thereby optimizing embeddings efficiently. This transformation is mathematically expressed as:
$\boldsymbol{x}_{\textrm{BSID}}=\boldsymbol{W}_{\textrm{emb}}[\textrm{BSID}].$ (1)
This embedding process, which involves learning the values of $\boldsymbol{W}_{\textrm{emb}}$ during training, effectively reduces both sparsity and dimensionality, thus simplifying the model and mitigating its computational demand. Additionally, the embedding allows the model to capture more complex relationships within the categorical BSID.
(a) Deep learning model
(b) Training and inference phase
Figure 2: Model Architecture: Our model is a two-layer Feed Forward Network. In the training phase, random masking is applied to randomly reassign BSIDs to ‘Unknown BS’.
Attention mechanism
To further enhance feature selection and to more effectively identify key features in the model, we propose an optimized network structure, incorporating a novel attention mechanism module, inspired by attention mechanisms utilized in Natural Language Processing (NLP) and Computer Vision (CV) applications. Our module, named adaptive re-weighting layer (ARL), enhances model predictions by assigning greater weights to the most significant input features. Our ARL module operates by initially compressing the dimensionality of the neural network input, $x$, using a linear layer. This compression is followed by a subsequent expansion in dimensionality. To facilitate non-linear processing, a rectified linear unit (ReLU) function is incorporated at the intermediate stage. The process completes with the application of a sigmoid activation function, yielding a vector that matches the input’s dimensionality. This vector comprises values in the range of 0 to 1, effectively redistributing weights across the input features. This mechanism is encapsulated in the following equation:
$\boldsymbol{w}_{\textrm{Att}}(\boldsymbol{x})=\text{Sigmoid}(\text{Linear}(\text{ReLU}(\text{Linear}(\boldsymbol{x})))).$ (2)
The attention vector, $\boldsymbol{w}_{\textrm{Att}}(x)$, is then applied to the original input through element-wise multiplication to obtain the re-weighted input of the network, $\boldsymbol{x}{{}^{\prime}}$, as shown below:
$\boldsymbol{x}{{}^{\prime}}=\boldsymbol{w}_{\textrm{Att}}(\boldsymbol{x})\cdot\boldsymbol{x}.$ (3)
This process ensures that the network assigns higher significance to the most relevant features, thereby enhancing its precision and effectiveness.
IV-C Training with Random BSID Masking
To enhance our model’s generalization across unknown BSs, we have integrated a distinct BSID category for such BSs. As illustrated in Fig. 2(b), during the training phase, 30% of the samples in each epoch are randomly assigned this ‘unknown BS’ identifier. This technique is designed to enable the model to handle unknown BSIDs during the inference phase, enhancing its ability to generalize and avoid overfitting to BSs in the test set.
Our model’s training objective is to minimize the MAPE, which is defined as follows.
$\textrm{MAPE}=\frac{\sum_{i=1}^{n}\left|{y_{i}}-\widehat{y}_{i}\right|}{\sum_{i=1}^{n}\left|{y_{i}}\right|},$ (4)
where $y_{i}$ represents the i-th actual energy consumption sample, and $\widehat{y}_{i}$ denotes the model’s corresponding estimation.
Training was conducted over 1000 epochs with a batch size of 512. Additionally, an early stopping mechanism was adopted, halting training based on the validation set performance to further prevent overfitting.
V Experiment and Analysis
In this section, we present the performance achieved by our model and a comprehensive analysis of experiments conducted to benchmark the performance of our model against a range of alternative methods.
V-A Overall performance
Our model underwent extensive training, spanning over 1000 epochs, on an RTX 3090 GPU, with each epoch averaging a duration of one second. The trained model was tested to estimate the energy consumption of the 10,076 samples included in the testing set, where it achieved remarkable performance, evidenced by a MAPE value of 4.9811%.
V-B BSID Generalization Capabilities
In this section, we discuss the strategies employed to enhance our model’s ability to generalize across different BSs by varying the processing of the BSID. We examined the impact of utilizing either One-Hot encoded BSID and embedding BSID or input features on model performance.
As indicated in Table II, employing One-Hot BSID in our model led to substantial predictive errors for unknown BSs. Without masked training, the model’s errors were more pronounced and performance worse than omitting BSID. However, using masked training significantly improved generalization for unknown BSs while maintaining accuracy for known BSs. Furthermore, the adoption of BSID embedding, in conjunction with masked training, led to a significant accuracy boost for unknown BSs, outperforming the results attained with One-Hot encoding. We also conducted experiments using the method proposed by [14] as a baseline. Due to its omission of the BSID as an input feature and a different encoding approach, it exhibited poor performance on this dataset.
To understand these dynamics, we employed Uniform Manifold Approximation and Projection (UMAP) for dimensionality reduction, compressing each BS’s embedding into a 2-dimensional space for visual analysis. As illustrated in Fig. 3, the 2-dimension visualization of BSID embeddings is color-coded according to RUType to demonstrate that it successfully captures the relationships between different BSs. This visualization reveals clustering patterns and proximity in the embeddings space between different RUTypes, showing that BSs of the same RUType are closer to each other in this space.
Figure 3: 2-Dimension visualization of BSID embeddings via UMAP. Different colors represent different RUTypes.
Fig. 4(a) and 4(b) show the energy consumption predictions outputted by our model for a specific known RUType, comparing scenarios where BSID embedding is either included or omitted. When the BSID is excluded, the model converges towards a generalized pattern, aligning energy consumption estimates with the average behavior observed across all BSs for each RUType. In contrast, the integration of BSID embeddings enables the model to capture and incorporate the distinctive attributes of each individual BS thereby refining the precision of its energy consumption predictions.
(a) Without BSID
(b) With BSID embedding
(c) Different encoding methods for ABF
Figure 4: Comparative visualization of predictions for a known RUType dataset: (a) Model without BSID, (b) Model with embedding BSID, and (c) One-Hot vs. Numerical encoding for ABF on an unknown RUType dataset.
TABLE II: Comparison of BSID encoding methods, assessing the impact of random masking on model parameters and MAPE (%). Best performances are highlighted in bold. RM is used to denote Random Masking.
Methods (Dim) Parameters Unknown-BS Known-BS All-BS Embedding BSID w/ RM(204) (Ours) 105209 6.6863 4.0 4.9811 Embedding BSID w/o RM(204) 105209 21.8093 3.9984 11.4294 One-Hot BSID w/ RM(1064) 286219 9.44 3.9452 6.1313 One-Hot BSID w/o RM(1064) 286219 20.9739 4.0732 11.2645 No BSID(140) 28757 8.1589 9.2191 8.7538 Baseline[14](55) 2871 13.9083 11.3856 12.7587
V-C Performance Comparison of One-Hot Encoding Combinations
In this section, we analyze the performance of our model in relation to different methods for incorporating some of the numerical features. Despite the inherent numerical nature of frequency, bandwidth, and the number of antennas, a consistent and clear linear relationship between these features and energy consumption is not always evident across all BSs, which may influence the precision of our model.
To address this issue, we explored treating these features both as numerical and as categorical variables. This approach entailed conducting a series of comparative analyses, varying the feature configurations between purely numerical and categorically encoded using One-Hot encoding. The primary aim of our experiments was to identify the configuration that achieves the lowest MAPE. The results of these investigations are detailed in Table III. In this table, “A”, “B”, and “F” denote Antennas, Bandwidth, and Frequency respectively. The configurations such as “ABF”, “BF”, “AF”, “AB”, “A”, “B”, and “F” suggest that these features were treated as categorical and preprocessed using One-Hot encoding. Notably, the “ABF” configuration emerges as the most effective, achieving the lowest MAPE for both known and unknown BSs.
Furthermore, Fig. 4(c) shows a visualization of energy consumption predictions, comparing the outcomes when the number of antennas, bandwidth, and frequency are processed as numerical features versus One-Hot features for a new RUType. This comparison illustrates that treating these inputs as categorical via One-Hot encoding yields energy consumption estimates that more accurately align with the groundtruth.
TABLE III: Performance comparison of different One-Hot encoding combinations on MAPE (%).
Combinations (Dim) Unknown-BS Known-BS All-BS
ABF (204) (Ours) 6.6863 4.0 4.9811
AB (172) 15.3823 4.5665 9.4701
AF (188) 13.0811 4.0847 7.9097
BF (199) 11.7292 4.1652 7.34
A (156) 20.094 4.7138 11.3221
B (167) 12.7202 4.2781 7.3687
F (183) 11.1229 4.3239 7.2736
Numerical (151) 13.7479 4.0733 8.3435
V-D Evaluating the Effectiveness of Attention Mechanism
In this section, we assess the efficacy of our proposed ARL through a rigorous ablation study. We compared the performance of two network configurations: one incorporating the ARL and the other devoid of it. The results, as delineated in Table IV, reveal a marked improvement in MAPE across all BS when the ARL is used, at the cost of a slight increase in the model’s parameters. Importantly, integration of ARL not only reduced the MAPE for known BSs but also significantly for unknown BSs, reducing their error from 8.5095% to 6.6863%, thus highlighting its ability to improve network generalization in various scenarios.
To evaluate if such performance improvement is due to the increased number of parameters, we introduce a model variant with three hidden layers (256-128-64) that facilitate a balanced comparison. This modification aimed to establish a baseline for comparison that is fair in terms of network complexity and capability. However, this comparison shows that a mere increase in parameters does not correlate with a significant enhancement in model performance.
TABLE IV: Performance comparison of models with and without attention mechanisms on MAPE(%).
Model Parameters Unknown-BS Known-BS All-BS w/ ARL (Ours) 105209 6.6863 4.0 4.9811 w/o ARL 100097 8.5095 4.0561 5.6571 3-hidden-layer w/o ARL 159233 9.0253 4.0815 5.959
VI Conclusion
In this paper, we presented an innovative deep learning model for estimating the energy consumption of 5G BSs, a critical component of the rapidly evolving 5G networks.
Different from existing methods, our method integrates BSID into the input features to capture different energy fingerprints across various BSs which enables for improved estimation accuracy. This integration is facilitated through an embedding layer, offering a more precise BSID representation with significantly reduced feature dimensions in comparison to One-Hot encoding. Moreover, the introduction of a masked training approach, combined with an ARL attention mechanism, has substantially enhanced the model’s generalization capabilities and accuracy.